OpenStudy (sw050399):

You can't factor out a variable from an integral and still have it be the same as it was originally right?

OpenStudy (anonymous):

you can if your not integrating with respect to said variable

OpenStudy (sw050399):

But you can't if you are integrating with respect to that variable, right?

OpenStudy (anonymous):

correct

jimthompson5910 (jim_thompson5910):

Yes something like \[\large \int x*f(x)dx = x*\int f(x)dx\] is NOT true, but something like \[\large \int y*f(x)dx = y*\int f(x)dx\] is true because y is considered a constant in this case

OpenStudy (sw050399):

Thank you! That's what I thought, but Imjust needed some reassurance. Thanks again!

jimthompson5910 (jim_thompson5910):

yw